The yp value of (D²− 3D+ 2)y = e^{3x}, choose the correct answer


A) 1/2 e^{3x}


B) e^{3x}


C) 2e^{3x}. D) 3e^{3x}


1
Expert's answer
2021-09-04T01:00:59-0400

"\left(D^2-3D+2\right)y=e^{3x}\\
\mathrm{A\:linear\:non-homogeneous\:ODE\:with\:constant\:coefficients\:has\:the\:form\:of}\:\left(a_nD^n+...+a_1D+a_0\right)y=g\left(x\right)\\
\mathrm{The\:general\:solution\:to}\:\left(a_nD^n+...+a_1D+a_0\right)y=g\left(x\right)\:\mathrm{can\:be\:written\:as}\\
y=y_h+y_p\\
_h\mathrm{\:is\:the\:solution\:to\:the\:homogeneous\:ODE}\:\left(a_nD^n+...+a_1D+a_0\right)y=0\\
y_p\mathrm{,\:the\:particular\:solution,\:is\:any\:function\:that\:satisfies\:the\:non-homogeneous\:equation}\\
y=c_1e^{2x}+c_2e^x\\
\mathrm{The\:general\:solution\:}y=y_h+y_p\mathrm{\:is:}\\
y=c_1e^{2x}+c_2e^x+\frac{1}{2}e^{3x}"